OpenStudy (anonymous):

please help on solving x^2+5x+3=0 and During the first part of a canoe trip, Ken covered 60km at a certain speed. He then traveled 24km at a speed that was 4 km/h slower. If the total time for the trip was 8 hr, what was the speed on each part of the trip? Thanks

hero (hero):

Would've been better to post these one question at t atime

OpenStudy (anonymous):

ok thanks ill fix it this is only the second time i use this help website because i'm pretty good at math but this subject is confusing for me.

hero (hero):

\[(x + \frac{5+\sqrt{13}}{2}) (x + \frac{5-\sqrt{13}}{2}) \]

OpenStudy (anonymous):

where did you get the 13 from?

OpenStudy (turingtest):

x^2+5x+3=0 does not factor, so you must use the quadratic formula:\[x={-b \pm \sqrt{b^2-4ac}\over2a}\]for equations of the form ax^+bx+c=0 here x^2+5x+3=0 a=1 b=5 c=3 plug in and solve!

hero (hero):

Turing Test, would've been nice if you posted your solution

hero (hero):